I drive 60 miles in 48 minutes and 10 miles at 35 miles per hour. What is my average speed and what is my average acceleration?



Answer by stanbon(75887)

I drive 60 miles in 48 minutes and 10 miles at 35 miles per hour. What is my average speed and what is my average acceleration?
---------------
To get average speed:
time = d/r = 10 mi/35 mph = 2/7 hrs
---
Average speed = total distance/total time = (60 + 10)mi/((48/60)+(2/7)) hrs
= 70/1.0857 = 64.47 mph
---------------------------------
To get average acceleration:
64.47/1.0857 = 59.38 mi/hr^2
====================================
Cheers,
Stan H.